8. A friend asks to borrow $49 from you and in return will pay you $52 in one year. If your bank is offering a 5.9% interest rate on deposits and​ loans:

a. How much would you have in one year if you deposited the $49 ​instead?

b. How much money could you borrow today if you pay the bank $52 in one​ year?

c. Should you loan the money to your friend or deposit it in the​ bank?
Answer #1

Answer:

a. Amount in 1 year = 49 * 1.059 = 51.891

b. amount to be borrowed today = 52/1.059 = 49.1029

c. since the bank is giving 51.891 it is better to deposit in the bank.